OpenStudy (anonymous):

Which of the following is the largest volume? 1.2 x 105 cm3 1.2 x 102 m3 2.0 x 104 dm3 1.2 x 108 mm

OpenStudy (anonymous):

1.2 x 102 m3

OpenStudy (anonymous):

thanks thats what i figured out also!

thomaster (thomaster):

Best way to do this is convert them all to the same volume unit, like liters Then it would be: \(\sf\large 1.2 x 10^5 cm3= 120 L \) \(\sf\large 1.2 x 10^2 ~~m3 =120,000 L\) \(\sf\large 2.0 x 10^4 dm3 =20,000 L\) \(\sf\large 1.2 x 10^8 mm \) (millimeter) is not a volume unit so it can't be converted.